Show that {fn}, where fn : [0, 1] → R is the map fn(x) = x
n, is not uniformly convergent.
1
Expert's answer
2020-06-22T17:32:05-0400

Given "f_n: \ [0,1] \to R" such that "f_n(x) =x^n" .

Now, Pointwise convergence is "\lim_{n\to \infin} f_n(x) = \lim_{n\to \infin} x^n = \begin{cases} 0 \ if \ x\in[0,1) \\ 1 \ if \ x=1 \end{cases}" .

Each "f_n(x) =x^n" is continuous for every n but it's pointwise limit is not continuous. So "f_n(x)" is not uniformly convergent.